# Split the Yield, Share the Risk: Pricing, Hedging and Fixed rates in DeFi

### v2

@@ 5 Fixed interest rate lending
Each bonding curve in the menu presented in Corollary [4.9] can now be closely approximated by a concentrated liquidity position [76] in a canonical Uniswap-v3 style market maker [71]. To see this, note that the liquidity in an infinitesimally narrow price range $[p,p+dp]$ of the curve corresponding to the $t^{th}$ yield future can be approximated by the expression $\frac{dx}{d(\sqrt{p})}$ [71], where $x$ is the amount of numeraire in the AMM reserve that obeys the LP’s bonding curve $\psi$. One can then convert all bonding curves in Corollary [4.9] to a liquidity distribution over price ranges (Detailed construction in Appendix [G]). This liquidity distribution is additive across all LPs, thus providing a unified (hence, more liquid) market for not only yield tokens of the same maturity and different price preferences, but also yield tokens of different maturities. As in the canonical Uniswap v3 liquidity provision, the LP decides which price ranges (equivalent to an implied yield range as we saw in Section 2.2) to provide liquidity. After that, the amount of yield tokens and the price range determines the liquidity level of the bonding curves and hence the amount of numeraire required to be deposited.
@@
Apart from bringing more liquidity into lending pools, using a yield token AMM acts as a lending liquidity aggregator by harnessing the DEX aggregation capabilities that already exist. Thus, loans of every maturity can be viewed as an orderbook [45] with a specific market price. Comparing these interest rates across varying collateral, any user would be able to lend/borrow tokens at the best possible interest rate, with the depth of the orderbook serving as a proxy for how much the market is certain about the current interest rates. We give explicit details of the procedure outlined above to convert the bonding curves from Corollary [4.9] to an aggregated order book in Appendix [G].
@@ 6 Conclusion and Future Work
Tokenizing yield for LP tokens, liquid staking tokens also opens up hedging opportunities for users. In the case of liquid staking tokens especially, yield tokenization can lead to gas fee hedging and slashing risk insurance. We provide some preliminary results in Appendix [H], with the full design of slashing insurance as an open question. On the flip side, a liquid yield market for staking may exacerbate the principal-agent problems in liquid staking/restaking (See for e.g. [74] and Appendix [H]). We leave the fair pricing and quantifying the systemic risk and/or consensus related externalities of other such yield tokens to future work.
Appendix A Proof of Corollary [2.9]
Proof A.1.
@@
We can now form a portfolio consisting of $\frac{\partial P_{\mathcal{Y}}^{T_{2}}}{\partial X}$ amount of the yield token with maturity $T_{1}$ and $-\frac{\partial P_{\mathcal{Y}}^{T_{1}}}{\partial X}$ amount of the yield token with maturity $T_{2}$. Doing this cancels out the $dW$ and gives only the risk free drift term $dt$. The coefficient of $dt$ is nothing but the PDE derived above (Equation [12]), of which the fair prices in Theorem 2.7 are a solution. This implies that the drift term must be zero. This implies that there is no risk free arbitrage opportunity between a pair of yield tokens with different maturities.
Appendix B Proof of Theorem 3.5
Proof B.1.
@@
    Assume that the lending pool has a utilization $u\in(0,1]$. Then, if the total amount of the lent asset in the pool is $L=\sum L_{i}$, the total amount being borrowed is $uL=\sum B_{i}$. From the proof of Lemma 3.3, we see that, at the fair price $E[Y]$, all lenders (no matter what level of risk aversion) are willing to sell their yield tokens and all borrowers (no matter what level of risk aversion) are willing to buy yield tokens. The number of yield tokens created by the lenders would be $LY/Y=L$, while the number of yield tokens required by borrowers to completely hedge their positions is also $(\sum B_{i})Y/u=L$. This proves the first statement.
@@
Appendix C Proof of Lemma 4.1
Proof C.1.
@@
Comparing this with Equation [29] implies that $p_{1}>p_{2}$.
Appendix D Proof of Theorem 4.3
Proof D.1.
From the proof of Lemma 4.1, we see that $p_{\Delta}^{S}$ is exactly how we defined $p_{i}$ corresponding to a $\Delta_{i}$. This implies that the liquidity provider faces the same expected utility over her wealth at maturity no matter where her reserves move along the curve $\psi_{S}$ as defined. A similar argument holds for the buy side of the curve using Lemma 4.2.
By the above lemmas, we also see that $p^{S}(\Delta)$ and $p^{B}(\Delta)$ are non-increasing and non-decreasing functions of $\Delta$ respectively. Therefore, the curve $\psi$ that is parametrized by $\Delta$ is convex and non-increasing when written as a function of $x$ in the $xy-$space. This implies that this curve is a well-defined, incentive compatible AMM bonding curve [3].
Appendix E Proof of Lemma 4.6
Proof E.1.
@@
by definition. Since $U_{S}$ is an increasing function, this equation can be satisfied by a single unique value of $p_{t}^{*}(\Delta)$.
Appendix F Proof of Lemma 4.7
Proof F.1.
@@
One can now define a new menu $\{p^{\prime}_{i}\}_{i=1}^{N}$ such that $p^{\prime}_{i}=p^{*}_{i}$ for $i\neq t^{*}$ and $p^{\prime}_{t^{*}}=p^{*}_{t^{*}}-\delta$, where $\delta>0$ is still small enough so that the optimal action of the trader is to sell the yield token at time $t^{*}$ over selling at $t$. However, for this new menu, we will have the utility of the LP to be $U_{P}^{\prime}:=\operatorname{\mathbb{E}}[U_{P}(x-y(\sum_{i=1}^{n}Y_{i})-\Delta(p^{\prime}_{t^{*}})+\Delta(\sum_{i=t^{*}}^{n}Y_{i}))]$. Since $U_{P}$ is an increasing function, we have that the value $U_{P}^{\prime}$ must be greater than the value $U_{P}^{*}:=\operatorname{\mathbb{E}}[U_{P}(x-y(\sum_{i=1}^{n}Y_{i})-\Delta(p^{*}_{t^{*}})+\Delta(\sum_{i=t^{*}}^{n}Y_{i}))]$ that the LP gets when presenting the menu $\{p_{i}^{*}\}_{i=1}^{N}$. This contradicts our assumption the $\{p_{i}^{*}\}_{i=1}^{N}$ is utility maximizing for the LP.
Appendix G Yield AMM aggregation
@@
In contrast, Uniswap v3’s concentrated liquidity model allows LPs to specify a price range $[p_{l},p_{u}]$ within which they provide liquidity. The amount of liquidity $L$ provided within this range determines the virtual reserves of the assets at any price $p$ within the range:
$x(p)=L(p^{\theta-1}-p_{u}^{\theta-1}),\quad y(p)=L(p^{\theta}-p_{l}^{\theta})$
@@
[…]
@@
Appendix H Yield tokenization in liquid staking
@@
H.1 Fair price of risk-free staking yield
@@
Lemma H.1.
Lemma H.3 Assume that the liquid staking token $\mathcal{A}$ has been locked in the yield tokenizer for only one time slot with yield at maturity $Y$. Then, the fair prices of the principal token and yield token are given by
@@
This implies that the fair price for the principal token must be as stated.
@@ H.2 Externalities - Insurance
Proof H.5.
Let $K$ denote the number of principal tokens bought by the liquid staker. Using Lemma H.3, the expected profit from purchasing the insurance is
@@
Here, the first term represents the case when no slashing occurs, in which the staker loses the insurance premium (equal to the price of the extra principal tokens that were bought). The second term represents the profit obtained from selling the principal tokens in case of slashing. The last term arise from selling the original principal token that the staker minted for free for depositing her liquid staking token in the yield tokenizer. For fairness, the expected loss to insurance must be 00. Setting the above expression to 00 and solving for $K$ gives us the desired result.

### v3

@@ 5 Fixed interest rate lending
Each bonding curve in the menu presented in Corollary [5] can now be closely approximated by a concentrated liquidity position [76] in a canonical Uniswap-v3 style market maker [71]. To see this, note that the liquidity in an infinitesimally narrow price range $[p,p+dp]$ of the curve corresponding to the $t^{th}$ yield future can be approximated by the expression $\frac{dx}{d(\sqrt{p})}$ [71], where $x$ is the amount of numeraire in the AMM reserve that obeys the LP’s bonding curve $\psi$. One can then convert all bonding curves in Corollary [5] to a liquidity distribution over price ranges (Detailed construction in Appendix [J]). This liquidity distribution is additive across all LPs, thus providing a unified (hence, more liquid) market for not only yield tokens of the same maturity and different price preferences, but also yield tokens of different maturities. As in the canonical Uniswap v3 liquidity provision, the LP decides which price ranges (equivalent to an implied yield range as we saw in Section 2.2) to provide liquidity. After that, the amount of yield tokens and the price range determines the liquidity level of the bonding curves and hence the amount of numeraire required to be deposited.
@@
Apart from bringing more liquidity into lending pools, using a yield token AMM acts as a lending liquidity aggregator by harnessing the DEX aggregation capabilities that already exist. Thus, loans of every maturity can be viewed as an orderbook [45] with a specific market price. Comparing these interest rates across varying collateral, any user would be able to lend/borrow tokens at the best possible interest rate, with the depth of the orderbook serving as a proxy for how much the market is certain about the current interest rates. We give explicit details of the procedure outlined above to convert the bonding curves from Corollary [5] to an aggregated order book in Appendix [J].
@@ 6 Conclusion and Future Work
Tokenizing yield for LP tokens, liquid staking tokens also opens up hedging opportunities for users. In the case of liquid staking tokens especially, yield tokenization can lead to gas fee hedging and slashing risk insurance. We provide some preliminary results in Appendix [K], with the full design of slashing insurance as an open question. On the flip side, a liquid yield market for staking may exacerbate the principal-agent problems in liquid staking/restaking (See for e.g. [74] and Appendix [K]). We leave the fair pricing and quantifying the systemic risk and/or consensus related externalities of other such yield tokens to future work.
7 Acknowledgements
The authors thank a generous gift from XinFin Private Ltd for supporting this research. We also thank Xuechao Wang, Tarun Chitra, Ciamac Moallemi, Aviv Zohar, Nico Pei for fruitful comments and discussion.
Appendix A Proof of Theorem 1
Proof.
By Ito’s formula, we get
Suppose now, that we have a portfolio consisting of one unit of the yield token $\mathcal{Y}$ and $\frac{\partial P_{\mathcal{Y}}^{T}}{\partial X^{i}_{t}}$ amounts of the token $\mathcal{A}_{i}$, for all underlying tokens $\{\mathcal{A}_{i}\}$. Let the value of the portfolio in terms of the numeraire $\mathcal{N}$ be $V_{t}$. After a time $dt$, the change in the portfolio value is given by
Substituting Equation [15] in Equation [17] gives us
Also, we know that the value of the investment at time $t$ is given by
$V_{t}$ $=P_{\mathcal{Y}}^{T}-\nabla P_{\mathcal{Y}}^{T}\cdot\bar{X_{t}}$ (21)
where we have $P_{\mathcal{Y}}^{T}(T,x)=0$ as a boundary condition.
To solve this PDE, we first define
$Z(s)$ $:=e^{-\int_{t}^{s}r(u)du}P_{\mathcal{Y}}^{T}(s,\bar{X_{s}})$ (24)
$d\bar{X_{t}}=r(t)\bar{X_{t}}dt+\Sigma_{t}d\bar{W_{t}^{*}}$ (25)
$dZ_{s}$ $=e^{-\int_{t}^{s}r(u)du}\left(-r(s)P_{\mathcal{Y}}^{T}+\frac{\partial P_{\mathcal{Y}}^{T}}{\partial s}+\frac{1}{2}\nabla\cdot\left(\Sigma_{s}\Sigma_{s}^{T}\right)\nabla P_{\mathcal{Y}}^{T}+\bar{X_{s}}\cdot r(t)\nabla P_{\mathcal{Y}}^{T}\right)dt$ (26)
$+e^{-\int_{t}^{s}r(u)du}\Sigma_{s}d\bar{W}_{s}^{*}$ (27)
Taking expectations conditioned on $\bar{X_{t}}=x$ on both sides, using Equation [23] and integrating $s$ from $t$ to $T$, using the fact that the price of the yield token is 0 at maturity, gives us the result. ∎
Appendix B Proof of Corollary [1]
Proof.
@@ Appendix B Proof of Corollary [1]
We can now form a portfolio consisting of $\frac{\partial P_{\mathcal{Y}}^{T_{2}}}{\partial X}$ amount of the yield token with maturity $T_{1}$ and $\frac{\partial P_{\mathcal{Y}}^{T_{1}}}{\partial X}$ amount of the yield token with maturity $T_{2}$. Doing this cancels out the $dW$ and gives only the risk free drift term $dt$. The coefficient of $dt$ is nothing but the PDE derived above (Equation [23]), of which the fair prices in Theorem 1 are a solution. This implies that the drift term must be zero. This implies that there is no risk free arbitrage opportunity between a pair of yield tokens with different maturities. ∎
Appendix C Proof of Lemma [1]
Proof.
If the proposed equation does not hold, we have the following cases.
Appendix D Proof of Lemma [2]
Proof.
We have the following two cases.
Taking $\delta_{\mathcal{B}}$$\delta_{\mathcal{L}}$ as defined above, by Jensen’s inequality, we know that, for any $P_{\mathcal{Y}}^{T}\in[E[Y]-\delta_{\mathcal{L}},E[Y]+\delta_{\mathcal{B}}]$, we have $U_{\mathcal{B}}(-B-B(P_{\mathcal{Y}}^{T})\geq E[U_{\mathcal{B}}(-B-BY/u)]$ and $U_{\mathcal{L}}(W+W(P_{\mathcal{Y}}^{T})\geq E[U_{\mathcal{L}}(L+LY)]$. Thus, we see that a lender prefers to sell her yield tokens, while borrower prefers to buy yield tokens to hedge the interest rate payment on his loan. ∎
Appendix E Proof of Theorem 2
Proof.
@@ Appendix E Proof of Theorem 2
    Assume that the lending pool has a utilization $u\in(0,1]$. Then, if the total amount of the lent asset in the pool is $L=\sum L_{i}$, the total amount being borrowed is $uL=\sum B_{i}$. From the proof of Lemma [2], we see that, at the fair price $E[Y]$, all lenders (no matter what level of risk aversion) are willing to sell their yield tokens and all borrowers (no matter what level of risk aversion) are willing to buy yield tokens. The number of yield tokens created by the lenders would be $LY/Y=L$, while the number of yield tokens required by borrowers to completely hedge their positions is also $(\sum B_{i})Y/u=L$. This proves the first statement.
@@
Appendix F Proof of Lemma [3]
Proof.
@@ Appendix F Proof of Lemma [3]
Comparing this with Equation [29] implies that $p_{1}>p_{2}$. ∎
Appendix G Proof of Theorem 3
Proof.
From the proof of Lemma [3], we see that $p_{\Delta}^{S}$ is exactly how we defined $p_{i}$ corresponding to a $\Delta_{i}$. This implies that the liquidity provider faces the same expected utility over her wealth at maturity no matter where her reserves move along the curve $\psi_{S}$ as defined. A similar argument holds for the buy side of the curve using Lemma [4].
By the above lemmas, we also see that $p^{S}(\Delta)$ and $p^{B}(\Delta)$ are non-increasing and non-decreasing functions of $\Delta$ respectively. Therefore, the curve $\psi$ that is parametrized by $\Delta$ is convex and non-increasing when written as a function of $x$ in the xyxy-space. This implies that this curve is a well-defined, incentive compatible AMM bonding curve [3]. ∎
Appendix H Proof of Lemma [5]
Proof.
@@ Appendix H Proof of Lemma [5]
by definition. Since $U_{S}$ is an increasing function, this equation can be satisfied by a single unique value of $p_{t}^{*}(\Delta)$. ∎
Appendix I Proof of Lemma [6]
Proof.
@@ Appendix I Proof of Lemma [6]
One can now define a new menu $\{p^{\prime}_{i}\}_{i=1}^{N}$ such that $p^{\prime}_{i}=p^{*}_{i}$ for $i\neq t^{*}$ and $p^{\prime}_{t^{*}}=p^{*}_{t^{*}}-\delta$, where $\delta>0$ is still small enough so that the optimal action of the trader is to sell the yield token at time $t^{*}$ over selling at $t$. However, for this new menu, we will have the utility of the LP to be $U_{P}^{\prime}:=\operatorname{\mathbb{E}}[U_{P}(x-y(\sum_{i=1}^{n}Y_{i})-\Delta(p^{\prime}_{t^{*}})+\Delta(\sum_{i=t^{*}}^{n}Y_{i}))]$. Since $U_{P}$ is an increasing function, we have that the value $U_{P}^{\prime}$ must be greater than the value $U_{P}^{*}:=\operatorname{\mathbb{E}}[U_{P}(x-y(\sum_{i=1}^{n}Y_{i})-\Delta(p^{*}_{t^{*}})+\Delta(\sum_{i=t^{*}}^{n}Y_{i}))]$ that the LP gets when presenting the menu $\{p_{i}^{*}\}_{i=1}^{N}$. This contradicts our assumption the $\{p_{i}^{*}\}_{i=1}^{N}$ is utility maximizing for the LP. ∎
Appendix J Yield AMM aggregation
@@
In contrast, Uniswap v3’s concentrated liquidity model allows LPs to specify a price range $p_{l},p_{u}]$ within which they provide liquidity. The amount of liquidity $L$ provided within this range determines the virtual reserves of the assets at any price $p$ within the range:
$x(p)=2L(1-\theta)(p^{\theta-1}-p_{u}^{\theta-1}),\quad y(p)=2L\theta(p^{\theta}-p_{l}^{\theta})$
@@
A demand curve represents the minimum amount of an asset the liquidity provider wants to hold at a given price. Thus, any bonding curve $\psi(x,y)=L$ can be converted into a demand curve by simply expressing the reserves of the asset as a function of price $x(p)$. The reserves of the numeraire are implied by the demand since they are tied with each other via the price of the asset. That is, we have $y(p)=-\int_{\infty}^{p}\pi x^{\prime}(\pi)d\pi$. This representation of the AMM now easily allows us to split and aggregate liquidity across multiple liquidity providers. The demand curves for each LP can be simply added to give us the demand curve of the market as a whole. Another representation that gives us a much more intuitive handle on the notion of where liquidity is present, is the derivative of the demand curve $L(p)=-x^{\prime}(p)$, which we call the liquidity curve. This is a measure of how much the liquidity provider is willing to sell when the marginal price of the asset changes from $p$ to $p+dp$. This is the “orderbook” view of the market that is also additive across liquidity providers. To summarize, the AMM has three equivalent representations - the bonding curve [3] (constrains the reserves of the AMM directly in $x,y$-space), the demand curve [58](represents the minimum amount of asset that the AMM wants to hold as inventory at a price $p$), and the liquidity curve (represents the infinitesimal/marginal amount of asset that the AMM wants to sell at a certain price $p$). For the AMM to be incentive compatible (which is necessary for the AMM to be non-trivial and a valid price oracle), the bonding curve function has to be quasiconcave [3]. This implies that the demand curve has to non increasing [58] and hence the liquidity curve has to be non-negative. Continuing our concentrated liquidity example from above, we see that the liquidity curve corresponding to a basic Uniswap v3 position is $L(p)=\frac{L}{2p^{3/2}}\mathbbm{1}_{p\in[p_{l},p_{u}]}$.
@@
Appendix K Yield tokenization in liquid staking
@@
K.1 Fair price of risk-free staking yield
@@
Lemma 7.
Lemma [8] Assume that the liquid staking token $\mathcal{A}$ has been locked in the yield tokenizer for only one time slot with yield at maturity $Y$. Then, the fair prices of the principal token and yield token are given by
@@
This implies that the fair price for the principal token must be as stated. ∎
@@
Proof.
Let $K$ denote the number of principal tokens bought by the liquid staker. Using Lemma [8], the expected profit from purchasing the insurance is
@@
Here, the first term represents the case when no slashing occurs, in which the staker loses the insurance premium (equal to the price of the extra principal tokens that were bought). The second term represents the profit obtained from selling the principal tokens in case of slashing. The last term arise from selling the original principal token that the staker minted for free for depositing her liquid staking token in the yield tokenizer. For fairness, the expected loss to insurance must be 0. Setting the above expression to 0 and solving for $K$ gives us the desired result. ∎